Suppose that Lil John Industries’ equity is currently selling for $40 per share and that 3.3 million shares are outstanding. The firm also has 63,000 bonds outstanding, which are selling at 102 percent of par. Assume Lil John was considering an active change to its capital structure so that the firm would have a (D/E) of 1.5. How much would the firm have to sell?
Answer #1

Current Equity ratio = E/D+E

=(3.3 million * 40) / (63000*102)+(3.3 million *40)

=132000000/6426000+132000000

=132000000/138426000

=95.36%

Debt ratio

=D/D+E

=6426000/138426000

=4.64%

The current DE Ratio is 0.0464/0.9536 = 0.04868, so lil jhon would be contemplating to increase DE ratio. Thus to berind debt ratio of 1.5/2.5 ie 0.6 Wou;d required issuing (0.6-0.04868)*(3.3million*40)+(63000*102) = 76317022.2$ of new debt and using this proceed to repurchase the stock